OpenStudy (anonymous):

For the function f(x)=-2sin(4x-pi), in naming the variables, would D=0?

zepdrix (zepdrix):

D being the horizontal shift? Like from this form?\[\Large\rm y=A \sin(B(x-C))+D\] Yes D=0 :)

OpenStudy (anonymous):

I also have: A=-2 B=4 C=pi?

zepdrix (zepdrix):

I mean vertical shift :3 blah sorry.

zepdrix (zepdrix):

Umm I guess it depends how you're defining your horizontal shift. The horizontal shift is with the coefficient on the x pulled out first. But yes I still think we call it C, because we define our horizontal shift to be C/B. Ya ya ya ok ok ok, ignore my rambling. Yessss those values look correct \c:/ good job!

OpenStudy (anonymous):

Okay!!! Thank you so much! Just wanted to double check.

OpenStudy (anonymous):

Expanding on this, I have: Amplitude=-2 Period=pi/2 Phase shift=pi/4 Right end point=3pi/4?